Derive the integrated law for zero and second order kinetics.
Answer #1

considering Reaction - - → product Rate love 9) zero for order - Reaction :- Rate (R) = K [A] > 0 R = k K=Rate constant. forconcentration of A. For Zero order Reaction- from equation ③ & - Ki-d{A at » - dA] = + d{A] = -kat om in o both sides d[A7= 1Co-kt) = for Zero ordez Reaction . TEN: $43kt For second order Reaction from equatica o & — K[A??= -dEA] at 1992 --Kat integr[A]~2 d[A] . Ik dt 1 xndx = fant n+1 TEų [A7-2+1 42+1 * . -K[27 alo 2) [429 o-k [-o] -> +inn 7 . +32 - 1 tokч - K+ ) – 4 = Coq kt — » Со = 4+ №4 Kt — co= q (It cokt) Cokt әре . 1+k+ . Так — «К++) co C = It cokt = for second order Rea